# Angular dependent measurement of electron-ion recombination in liquid argon for ionization calorimetry in the ICARUS liquid argon time projection chamber

### v1

@@ 2 Particle Track Energy Scale
In our calibration procedure, for the muon population we elect to only include depositions where the analytic Landau approximation is valid. This limits us to depositions near the mimumum-ionizing-particle (MIP) region, where recombination is close to linear. To provide information on recombination, we also include highly-ionizing-particle (HIP) depositions from protons near their Bragg peak. In this case, the distribution of energy loss is approximately a Landau-Vavilov distribution [22], although there is a perturbation to the shape from diffusion [23]. The MPV of the Landau-Vavilov distribution cannot be expressed analytically. We use the numerical computation provided by the ROOT VavilovAccurate routine555We have found in our simulation that in this region the MPV of the proton energy loss distribution is modeled well by the ROOT numerical computation, using the track pitch (not the track thickness) as the input to the Landau parameter () and value used in the VavilovAccurate routine. [25].
@@
| Transverse diffusion constant 666This value is taken from an analysis in ICARUS in preparation to be published. The preliminary value applied here will not change in the final result by an amount significant enough to impact the results of this paper. | |
@@ 4.1 Cosmic-Ray Muon Selection
Cosmic-ray muon tracks are required to cross the central cathode plane in either cryostat. The identification of a track in both TPCs on either side of the cathode allows the arrival time of the muon to be determined. This enables the track charge to be corrected for attenuation due to argon impurities. These “cathode-crossing tracks" are identified by Pandora. The muon is required to stop within a fiducial volume (see table 2), and is required to have a median in the last along the track greater than 75 kecm 777The cut value is expressed directly in terms of analog-digital counts (ADC) per centimeter. The cut is . The written value is converted to a number of electrons () applying an approximate gain of 75 ADC.. This cut increases the stopping track purity by requiring the measured charge near the endpoint to be consistent with the expectation from the muon Bragg peak. The region of the track where the cut is applied is not included in the calibration, and so the cut should not bias the measurement. This selection identifies 93 thousand muon candidates. A Monte Carlo simulation study (with cosmic-ray muon generation modeled by CORSIKA [32]) of this selection yields a 93% purity of tracks stopping within of the reconstructed endpoint. The remaining 7% consist almost entirely of cosmic-ray muons where the endpoint is mis-reconstructed, either due to track splitting or combining with a Michel electron. The dependence of energy loss on the muon momentum is small for the range of muon momenta used in this measurement, so this impurity has a small impact.
@@ 4.3 Charge Scale Corrections
We use the Integral charge method for calibrating the ICARUS TPC energy scale, but use the SummedADC method to diagnose the angular dependence. The ratio of the Integral to SummedADC charge reconstruction is taken as a charge scale correction factor as a function of the track angle 888This angle determines the shape of the ionization charge signal from a track, and therefore controls any orientation dependence in the charge reconstruction. (see figure 2). Figure 4 plots the correction factor in Monte Carlo simulation and data. A systematic uncertainty of 0.2% on the correction is assigned to cover the differences.

### v2

@@ 2 Particle Track Energy Scale
In our calibration procedure, for the muon population we elect to only include depositions where the analytic Landau approximation is valid. This limits us to depositions near the mimumum-ionizing-particle (MIP) region, where recombination is close to linear. To provide information on recombination, we also include highly-ionizing-particle (HIP) depositions from protons near their Bragg peak. In this case, the distribution of energy loss is approximately a Landau-Vavilov distribution [22], although there is a perturbation to the shape from diffusion [23]. The MPV of the Landau-Vavilov distribution cannot be expressed analytically. We use the numerical computation provided by the ROOT VavilovAccurate routine666We have found in our simulation that in this region the MPV of the proton energy loss distribution is modeled well by the ROOT numerical computation, using the track pitch (not the track thickness) as the input to the Landau parameter () and value used in the VavilovAccurate routine. [25].
@@
| Transverse diffusion constant 777This value is taken from an analysis in ICARUS in preparation to be published. The preliminary value applied here will not change in the final result by an amount significant enough to impact the results of this paper. | |
@@ 4.1 Cosmic-Ray Muon Selection
Cosmic-ray muon tracks are required to cross the central cathode plane in either cryostat. The identification of a track in both TPCs on either side of the cathode allows the arrival time of the muon to be determined. This enables the track charge to be corrected for attenuation due to argon impurities. These “cathode-crossing tracks" are identified by Pandora. The muon is required to stop within a fiducial volume (see table 2), and is required to have a median in the last along the track greater than 75 kecm 888The cut value is expressed directly in terms of analog-digital counts (ADC) per centimeter. The cut is . The written value is converted to a number of electrons () applying an approximate gain of 75 ADC.. This cut increases the stopping track purity by requiring the measured charge near the endpoint to be consistent with the expectation from the muon Bragg peak. The region of the track where the cut is applied is not included in the calibration, and so the cut should not bias the measurement. This selection identifies 93 thousand muon candidates. A Monte Carlo simulation study (with cosmic-ray muon generation modeled by CORSIKA [32]) of this selection yields a 93% purity of tracks stopping within of the reconstructed endpoint. The remaining 7% consist almost entirely of cosmic-ray muons where the endpoint is mis-reconstructed, either due to track splitting or combining with a Michel electron. The dependence of energy loss on the muon momentum is small for the range of muon momenta used in this measurement, so this impurity has a small impact.
@@ 4.3 Charge Scale Corrections
We use the Integral charge method for calibrating the ICARUS TPC energy scale, but use the SummedADC method to diagnose the angular dependence. The ratio of the Integral to SummedADC charge reconstruction is taken as a charge scale correction factor as a function of the track angle 999This angle determines the shape of the ionization charge signal from a track, and therefore controls any orientation dependence in the charge reconstruction. (see figure 2). Figure 4 plots the correction factor in Monte Carlo simulation and data. A systematic uncertainty of 0.2% on the correction is assigned to cover the differences.